A coil of wire has a resistance of 38.0 Ω at 25° C and 43.7 Ω at 54° C. What is the temperature coefficient of resistivity?
Answer #1

let the temperature coefficient of resistivity is a

for R1 = 38 Ohm

R2 = 43.7 Ohm

delta T = 54 - 25 = 29 degree C

for the resistance

R2 = R1 *(1 + a * delta T)

43.7 = 38 * (1 + a * 29)

a = 0.00517 1/degree C = 5.17 *10^-3 1/degree C

the temperature coefficient of resistivity is 5.17 *10-3 1/degree C